# Sign changes of Kloosterman sums with moduli having at most two prime factors

### v1

@@ Abstract.
We prove that the Kloosterman sum $\text{Kl}(1,q)$ changes sign infinitely many times, as $q\rightarrow+\infty$ with at most two prime factors. As a consequence, our result is unconditional compared with Drappeau and Maynard’s (Proc. Amer. Math. Soc., 2019), in which the existence of Laudau-Siegel zeros is required. Our arguments contain the Selberg sieve method, spectral theory and distribution of Kloosterman sums along with previous works by Fouvry, Matomäki, Michel, Sivak-Fischler, and Xi.
@@ 1.1. Background
Let $q$ be a fixed positive integer. For arbitrary integers $m$ and $n$, the classical Kloosterman sums are defined by
$S(m,n;q)=\sum_{a(q)\atop(a,q)=1}e\left(\frac{ma+n\overline{a}}{q}\right),$
where $e(x)=e(2\pi ix/q)$. Kloosterman sums have originated from Poincaré [12] and Kloosterman [8]. Scholars are concerned about whether there exists some specific explicit expression or asymptotic formula? Unfortunately, these results are beyond our capabilities for now.
Kloosterman sums are one of the central research interests in analytic number theory and related to numerous applications in Diophantine equations and automorphic forms. A well-known estimate from Weil [16] for the individual Kloosterman sum is
$\left|S(m,n;p)\right|\leq 2p^{1/2}(m,n,p)^{1/2},$
@@
$\left|S(m,n;q)\right|\leq q^{1/2}(m,n,q)^{1/2}\tau(q),$
and the factor $\tau(q)$ above can be modified to $2^{\omega(q)}$($\omega(q)$ represents the number of different prime factors of $q$) by Estermann [2].
@@
In this paper, we mainly concern about the sign changes Kloosterman sums. Fouvry and Michel [3][4] pioneering proved that
@@
Recently, Drappeau and Maynard [1] proved that under the condition of the existence of Landau-Siegel zero, 23 can be reduced to 2. Motivated by their previous works, we showed that 23 can be improved to 2, unconditionally.
Our new approach is somewhat similar to Maynard’s idea [11] of applying a variant of the Selberg sieve to improve Zhang’s breakthrough [20] in the Twin Prime Conjecture without changing the level of the Bombieri-Vinogradov Theorem.
@@ 1.2. Our result
$\bullet$ If not specified, we use $a\ (q)$ to represent $a(\bmod\ q)$ and $\overline{a}$ to represent the multiplicative inverse element modulo $q$.
@@
$\bullet \mathbb{C}$ represents the complex plane.
$\bullet$ $\Re s$ represents the real part of $s$.
$\bullet$ $\binom{a}{b}$ is binomial coefficient.
$\bullet$ $\int_{(c)}$ represents $\int_{c-i\infty}^{c+i\infty}$.
@@ 2. Preliminary
2.1. A new sieve weight
Define Selberg sieve weight $\left(\lambda_{d}\right)$ and state as
@@ 2.1. A new sieve weight
where $F$ is a fixed smooth function supported on $[0,1]$ and vanishes at 0 to a suitable order (related to $\kappa$ and $l$ below).
Our purpose is to show
where $c_{0}>0$ is a constant. It suffices to prove
$R^{\pm}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\frac{\left|S(1,1;n)\right|\pm S(1,1;n)}{\sqrt{n}}\left\{\rho-\left(\frac{\kappa}{2}\right)^{\omega(n)}\right\}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}>0,$
where $g(X)$ is a fixed smooth function supported on [1,2], $\Pi_{l}$ is the product of the first $l$ primes.
@@
(1) $\kappa=4,\ l=10,\ F(x)=x^{\kappa+l}.$
Comparing with [17] and [18], we replaced $d\mid n$ with $d\mid n\Pi_{l}$. This idea is essentially due to Selberg [13], and our purpose of doing this is to more freely select the coefficients of $R_{1}(X)$ and $R_{3}(X)$ in the following argument.
In reviewing Selberg’s method within the context of the twin prime conjecture, as outlined in [13], we encounter a sieve weight given by
$\left(\sum_{d\mid n(n+2)}\lambda_{d}\right)^{2}.$
However, in the specific problem we are addressing, the prime factorization of $n+2$ remains indeterminate. Consequently, we aim to substitute $n+2$ with a finite product of primes. Furthermore, examining the proof presented in Section 5 of [18], we observe that in the sieve weight, the oscillation amplitude of $\lambda_{d}$ becomes significant when $d$ has small prime factors. To handle this technical issue, we introduce $\Pi_{l}$ to amplify the magnitude of possible cancellation coming from $\lambda_{d}$.
Alternatively, if we persist with the original sieve weight summed over divisors $d$, potential improvements to our results might be achieved by improving the level of the Selberg sieve, particularly in the context of Lemma 3.2. Nevertheless, enhancing this level presents a formidable challenge for us at present. Fortunately, by adopting another new sieve weight, we can circumvent this challenge.
2.2. Outline of the proof
@@
$R_{1}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2},$
$R_{2}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\text{Kl}(1;n)\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2},$
$R_{3}(X):=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\frac{\kappa}{2}\right)^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}.$
We will discuss $R_{1}(X)$, $R_{2}(X)$ and $R_{3}(X)$ in Section 4, 5 and Section 6, respectively. The relevant conclusion of $R_{1}(X)$ can be referred to [10] and [17]. $R_{2}(X)$ can be derived from a B-V Theorem for Kloosterman sum, which can be found in [4], [15] and [17]. For $R_{3}(X)$, it can be handled in a similar way as in [18].
To prove the Theorem 1.2, we need the following conclusions.
@@
For any sufficiently large number $X$, there exists some constant $C_{1}>0$ such that
$R_{1}(X)\geq(1+o(1))C_{1}\frac{X}{\left(\log X\right)^{21}}.$
Specifically, we choose
$C_{1}=\left(4^{10}10!\binom{14}{10}\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 0.0142\tilde{g}(1),$
@@
For any sufficiently large number $X$, we have
$R_{2}(X)=o\left(\frac{X}{\left(\log X\right)^{21}}\right).$
Proposition 2.3.
For any sufficiently large number $X$, there exists some constant $C_{2}>0$ such that
$R_{3}(X)\leq C_{2}\frac{X}{\left(\log X\right)^{21}}(1+o(1)).$
Specifically, we choose
$C_{2}=4^{21}2^{10}\left(\prod_{p\mid\Pi_{10}}\log p\right)^{2}\times 8817.853\tilde{g}(1).$
Proposition 2.1, 2.2, 2.3 implies Theorem 1.2 immediately by choosing a suitable $\rho$ (for instance $\rho=5$, since $C_{1}>2C_{2}$) such that $52^{\omega(n)}<$ 5, which leads to $\omega(n)\leq 2$.
@@ 3. Some Lemmas
$H(X)=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\frac{\left|S(1,1;n)\right|}{\sqrt{n}}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
@@
$H(X)\geq\tilde{g}(1)\sum_{2\leq i\leq 5}2^{i}A_{i}(F)C_{i}\cdot\frac{X}{\log X}(1+o(1)),$
where $C_{2}=0.11109$, $C_{3}=0.03557$, $C_{4}=0.01184$ and $C_{5}=0.00396$,
$A_{i}(F)=\mathop{\int\dots\int}_{R_{i}}\frac{L_{i}^{2}\left(F;X^{1-\alpha_{2}-\dots-\alpha_{i}},X^{\alpha_{2}},\dots,X^{\alpha_{i}}\right)}{\alpha_{2}\dots\alpha_{i}(1-\alpha_{2}-\dots-\alpha_{i})}\text{d}\alpha_{2}\dots\text{d}\alpha_{i},$
$L_{i}(F;\alpha_{1},\alpha_{2},\dots,\alpha_{i})=\sum_{\mathcal{A}\subseteq\left\{\alpha_{1},\alpha_{2},\dots,\alpha_{i}\right\}\atop\sum_{\alpha\in\mathcal{A}}\alpha<\frac{1}{4}}(-1)^{|\mathcal{A}|}F\left(1-4\sum_{\alpha\in\mathcal{A}}\alpha\right),$
@@
$R_{2}:=$$\left\{\alpha_{2}\in[\eta,1):\left(\frac{3}{4}+\eta\right)(1-\alpha_{2})<\alpha_{2}<\frac{1}{2}\right\},$
@@
$\cap\left\{(\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5})\in[\eta,1)^{4}:\alpha_{5}<\alpha_{4}<\alpha_{3}<\alpha_{2}<1-\alpha_{2}-\alpha_{3}-\alpha_{4}\alpha_{5}\right\},$
$\eta:=$ $10^{-2023}.$
@@
Before presenting the following lemmas, let’s first establish some definitions. Define a smooth function $F(x)$ and assume it has a Taylor expansion
$F(x):=\sum_{n\geq 0}\frac{a_{n}}{n!}x^{n}$
with $F(0)=0$. For $N>1$ and $\0s\in\mathbb{C}\backslash\left\{0\right\}$, we further define
$\check{F}_{N}(s):=\sum_{n\geq 0}\frac{a_{n}}{(s\log N)^{n}}.$
@@
Let $N>1$ be not an integer. For any coefficient $y_{n}$ with $\ll\tau(n)^{O(1)}(\log n)^{O(1)}$, we have
$\sum_{n\leq N}y_{n}F\left(\frac{\log(N/n)}{\log N}\right)=\frac{1}{2\pi i}\int_{2-i\infty}^{2+i\infty}\check{F}_{N}(s)Y(s)\frac{N^{s}}{s}\text{d}s,$
@@
Let $v$, $v_{1}$, $v_{2}$ be fixed positive integers, $m$ be fixed even positive integer, $M>1$ be any real number. Suppose $P$, $Q$ are two smooth functions that have zeros of orders at least $v_{1}+m/2$, $v_{2}+m/2$ at 0, respectively. Let $Z(s_{1},s_{2})$ be holomorphic in the right half plane containing a neighborhood of $(0,0)$ with
@@
$R=(1+o(1))Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{\Gamma(v)m!}\int_{0}^{1}P^{(v_{1}+m/2)}(x)Q^{(v_{2}+m/2)}(x)(1-x)^{v-1}\text{d}x.$
@@
We follow the proof of Lemma A.2 in [18] and use Goldston et al.’s method in [5]. Assume
@@
From the condition of Lemma 3.4 this can be expanded as
@@
×$\times(\xi+1)^{v+k_{1}-v_{1}+k_{2}-v_{2}-i}(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-i}.$
We only need to consider the term with the quantity $(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-m}$, thus substituting into (3) to derive
$\frac{1}{(v+k_{1}-v_{1}+k_{2}-v_{2})!}\binom{v+k_{1}-v_{1}+k_{2}-v_{2}}{m}$ $Z^{(m)}(0,0)(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-m}$
(4) ×$\times\frac{1}{2\pi i}\int_{C_{3}}\frac{(\xi+1)^{k_{1}-v_{1}+k_{2}-v_{2}-m}}{\xi^{k_{2}-v_{2}-m/2+1}}\text{d}\xi.$
@@
$\begin{cases}&k_{1}-v_{1}\geq\frac{m}{2}\\ &k_{2}-v_{2}\geq\frac{m}{2},\end{cases}\ \begin{cases}&k_{1}-v_{1}<\frac{m}{2}\\ &k_{2}-v_{2}<\frac{m}{2},\end{cases}\ \begin{cases}&k_{2}-v_{2}\geq\frac{m}{2}\\ &k_{1}-v_{1}<\frac{m}{2},\end{cases}\ \begin{cases}&k_{2}-v_{2}>\frac{m}{2}\\ &k_{1}-v_{1}\geq\frac{m}{2}.\end{cases}$
The integral values for the last three cases are all equal to 0. When $k_{1}-v_{1}<\frac{m}{2}$ and $k_{2}-v_{2}<\frac{m}{2}$, only one pole $\xi=-1$ of order $m-(k_{1}-v_{1})-(k_{2}-v_{2})$ is in the region bounded by the circle $|\xi|=2$. Then Cauchy’s residue theorem implies that
@@
When $k_{1}-v_{1}<\frac{m}{2}$ and $k_{2}-v_{2}\geq\frac{m}{2}$, we divide this case into two cases, namely
@@
Only one pole $\xi=0$ of order $k_{2}-v_{2}-m/2+1$ is in the region bounded by $|\xi|=2$. Thus Cauchy’s residue theorem implies that
@@
and two poles $\xi=-1$ and $\xi=0$ are in the region bounded by $|\xi|=2$. Surprisingly, in this case, we can obtain
@@
thus the integral is still equal to 0.
The last case can be handled in a similar way to the third one. For the convenience of calculation, in this case, the positions of $s_{1}$ and $s_{2}$ are better to be interchanged, i.e., let $|s_{1}|=2\rho$, $|s_{2}|=\rho$.
Combining the above, only the case
@@
$(1+o(1))Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{m!}$
×$\ \ \ \ \ \ \ \times\sum_{k_{1}\geq v_{1}+m/2}\sum_{k_{2}\geq v_{2}+m/2}\frac{a_{k_{1}}b_{k_{2}}}{(v+k_{1}-v_{1}+k_{2}-v_{2}-m)!}\binom{k_{1}-v_{1}+k_{2}-v_{2}-m}{k_{2}-v_{2}-m/2}$
$=$ $(1+o(1))Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{m!\Gamma(v)}\int_{0}^{1}P^{(v_{1}+m/2)}(x)Q^{(v_{2}+m/2)}(x)(1-x)^{v-1}\text{d}x,$

### v2

# Sign changes of Kloosterman sums with moduli having at most six prime factors

@@ Abstract.
We prove that the Kloosterman sum $\text{Kl}(1,q)$ changes sign infinitely many times, as $q\rightarrow+\infty$ with at most six prime factors. As a consequence, our result improved the best known result of Xi(IMRN, 2022). The novelty of our method comes from introducing a new truncated divisor function whose selection depends on the number of prime factors of the variable, through which Kloosterman sum is controlled good enough. Our arguments contain the Selberg sieve method, spectral theory and distribution of Kloosterman sums along with previous nice works by Fouvry, Matomäki, Michel, Sivak-Fischler and Xi.
@@ 1.1. Background
Let $q$ be a fixed positive integer. For arbitrary integers $m$, the classical Kloosterman sums are defined by
$\text{Kl}(m;q)=\frac{1}{\sqrt{q}}\sum_{b(q)\atop(b,q)=1}e\left(\frac{mb+\overline{b}}{q}\right),$
where $e(x)=e^{2\pi ix}$. Kloosterman sums have originated from Poincaré [12] and Kloosterman [9]. Scholars are concerned about whether there exists some specific explicit expression or asymptotic formula? Unfortunately, these results are beyond our capabilities for now.
Kloosterman sums are one of the central topics in analytic number theory and related to numerous applications in Diophantine equations and automorphic forms. A well-known estimate from Weil [15] for the individual Kloosterman sum is
$\left|\text{Kl}(m;p)\right|\leq 2(m,p)^{1/2},$
@@
$\left|\text{Kl}(m;p)\right|\leq(m,q)^{1/2}\tau(q),$
and the factor $\tau(q)$ above can be modified to $2^{\omega(q)}$ for $32\nmid q$ by Estermann [2] ($\omega(q)$ represents the number of different prime factors of $q$).
@@
At present, we know very little about the “horizontal” Sato-Tate conjecture, but we still believe in its correctness. Part of the evidence is due to the numerical calculations of Kloosterman sums are highly consistent with “horizontal” Sato-Tate conjecture [11], and the other part is due to certain exponential sums such as cubic Gaussian sums and Salié sums do satisfy “horizontal” Sato-Tate conjecture. About the relevant applications and progress of Kloosterman sums can refer to [11].
In this paper, we mainly concern about a slightly weaker problem which is the sign changes of Kloosterman sums. Fouvry and Michel [3, 4] pioneeringly proved that
@@
Recently, Drappeau and Maynard [1] proved that under the condition of the existence of Landau-Siegel zero, 23 can be reduced to 2.
@@ 1.2. Our result
It seems that this conclusion is very difficult to be improved. The main reasons are to be stated in Section 7.
@@
$\bullet$ If not specified, we use a(q)a\ (q) to represent $a(\bmod\ q)$ and $\overline{a}$ to represent the multiplicative inverse modulo $q$.
@@
$\bullet$ $\binom{a}{b}$ is the binomial coefficient.
$\bullet$ $\widetilde{g}$ is the Mellin transform of $g$.
$\bullet$ $\left\{\mathcal{A}\right\}$ means the cardinality of the set $\mathcal{A}$.
$\bullet$ $\lVert\boldsymbol{a}\rVert$ represents $l_{2}$-norm.
$\bullet$ $\mathop{{\sum}^{*}}_{a(q)}$ represents the summation of $1\leq a\leq q$ with $(a,q)=1$.
@@ 2. Preliminary
2.1. Truncated divisor function
Define a new truncated divisor function
$\tau(n;\alpha,\beta)=\sum_{n=dl\atop d\in\mathcal{B}(n)}\alpha^{\omega(d)}\beta^{\omega(l)}$
with
(1) $\mathcal{B}(n)=\left\{d\in\mathbb{Z}^{+}:d\mid n,\omega(d)=3,\omega\left(\frac{n}{d}\right)\geq 3,d\leq n^{\frac{1}{2}}\right\}$
for square-free numbers $n$.
For our purposes, we need to select a subset from
$\left\{d\in\mathbb{Z}^{+}:d\mid n,d\leq n^{\frac{1}{2}}\right\}.$
We hope that the number of elements in this subset is appropriately small. After a long time of checking, we select (1) as the very required subset. What we need to emphasize is that this selection is quite different from the previous one in that it depends further on the number of prime factors of the variable, not just the size and range of the variable.
The reason for this selection will be stated in Section 7.
2.2. Outline of the proof
Our purpose is to show that
where $c_{0}>0$ is a constant. Following much the same way in [16, 17, 19], let $g(x)$ be a fixed smooth function supported on $[1,2]$, $\left(\lambda_{d}\right)$ be the Selberg sieve weight satisfying
@@ 2.2. Outline of the proof
where $F$ is a fixed smooth function supported on $[0,1]$ and vanishes at 0 to a suitable order (related to $k$ below, i.e. the dimension of the sieve problem). This choice of the order of $F$ is optimal, see [5] (Page 1090) for further discussion.
@@
$F(x)=x^{4}(a_{0}+a_{1}x+a_{2}x^{2}+a_{3}x^{3}+a_{4}x^{4}),$
with
$(a_{0},a_{1},a_{2},a_{3},a_{4})\approx\left(\frac{100396}{53901},-\frac{17284}{13475},\frac{76486}{134753},-\frac{33241}{188654},\frac{10836}{377308}\right).$
Define
(2) $R^{\pm}(X):=\sum_{\left(n,\Pi_{\varepsilon}\right)=1}g\left(\frac{n}{X}\right)\mu^{2}(n)\left(\left|\text{Kl}(1;n)\right|\pm\text{Kl}(1;n)\right)\left\{\rho-\tau\left(n;\alpha,\beta\right)\right\}\left(\sum_{d\mid n}\lambda_{d}\right)^{2},$
where $\rho$, $\alpha$, $\beta$ are fixed positive real numbers to be chosen later, $\Pi_{\varepsilon}=\prod_{p<X^{\varepsilon}}p$ is the prime factor product not exceeding $X^{\varepsilon}$ with a suitably small constant $\varepsilon>0$.
@@
$R_{1}(X):=\sum_{\left(n,\Pi_{\varepsilon}\right)=1}g\left(\frac{n}{X}\right)\mu^{2}(n)\left(\left|\text{Kl}(1;n)\right|\pm\text{Kl}(1;n)\right)\left(\sum_{d\mid n}\lambda_{d}\right)^{2},$
$R_{2}(X):=\sum_{\left(n,\Pi_{\varepsilon}\right)=1}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\tau\left(n;\alpha,\beta\right)\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
The lower bound for $R_{1}(X)$ can be divided into two parts, one of which can be obtained in [4], [10], [14] and [16], and the other part can be processed in a similar way as in [17].
For $R_{2}(X)$, we will further analyze this newly introduced divisor function. Part of the process can be handled in a similar way as in [19]. Building upon (1), we need to analyze some special forms of Lemma 3.3, which we will discuss in detail later.
To prove Theorem 1.2, we need the following conclusions.
@@
For any sufficiently large number $X$, we have
$R_{1}(X)\geq(1+o(1))0.76235\times\widetilde{g}(1)\frac{X}{\log X},$
@@
For any sufficiently large number $X$, suppose
$\alpha>0,\ \beta>0,$
and choose $\alpha=\alpha_{0}>0$, $\beta=1$, specifically. Then we have
$R_{2}(X)\leq(1+o(1))6.27044\alpha_{0}^{3}\times\widetilde{g}(1)\frac{X}{\log X}.$
@@ 3. Some Lemmas
$H(X)=\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)\left|\text{Kl}(1;n)\right|\left(\sum_{d\mid n}\lambda_{d}\right)^{2}.$
@@
$H(X)\geq\widetilde{g}(1)\sum_{2\leq i\leq 5}2^{i}C_{i}A_{i}(F)\cdot\frac{X}{\log X}(1+o(1)),$
where $C_{2}=0.11109$, $C_{3}=0.03557$, $C_{4}=0.01184$, $C_{5}=0.00396$,
$A_{i}(F)=\mathop{\int\ldots\int}_{R_{i}}\frac{L_{i}^{2}\left(F;X^{1-\alpha_{2}-\cdots-\alpha_{i}},X^{\alpha_{2}},\ldots,X^{\alpha_{i}}\right)}{\alpha_{2}\cdots\alpha_{i}(1-\alpha_{2}-\cdots-\alpha_{i})}\text{d}\alpha_{2}\cdots\text{d}\alpha_{i},$
$L_{i}(F;\alpha_{1},\alpha_{2},\ldots,\alpha_{i})=\sum_{\mathcal{A}\subseteq\left\{\alpha_{1},\alpha_{2},\ldots,\alpha_{i}\right\}\atop\sum_{\alpha\in\mathcal{A}}\alpha<\frac{1}{4}}(-1)^{|\mathcal{A}|}F\left(1-4\sum_{\alpha\in\mathcal{A}}\alpha\right),$
@@
$R_{2}:=$ $\left\{\alpha_{2}\in[\eta,1):\left(\frac{3}{4}+\eta\right)(1-\alpha_{2})<\alpha_{2}<\frac{1}{2}\right\},\ \eta:=10^{-2016},$
@@
$\cap\left\{(\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5})\in[\eta,1)^{4}:\alpha_{5}<\alpha_{4}<\alpha_{3}<\alpha_{2}<1-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}\right\}.$
@@
Before giving the following lemma, we first recall the “Siegel-Walfisz” condition. Let $\boldsymbol{v}=(v_{n})$ be an arithmetic function. For any positive integers $w$, $d$ with $(w,a)=1$, $a\neq 0$, if
$\sum_{n\leq X\atop{n\equiv a(w)\atop(n,d)=1}}v_{n}-\frac{1}{\phi(w)}\sum_{n\leq X\atop(n,dw)=1}v_{n}\ll\lVert\boldsymbol{v}\rVert X^{\frac{1}{2}}\tau(d)^{B}\mathcal{L}^{-A}$
holds for some constant $B>0$ and any $A>0$ with an implied constant in $\ll$ depending only on $A$, then we say $\boldsymbol{v}$ satisfies the “Siegel-Walfisz” condition [18].
@@
Let $\boldsymbol{\alpha}=(\alpha_{m})$ be a complex coefficient with support in $[M,2M]$ satisfying the above “Siegel-Walfisz” condition. For square-free $q\geq 1$, put
$\varepsilon(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}_{q};q)=\frac{1}{\sqrt{q}}\mathop{{\sum}^{*}}_{r(q)}\left|\text{Kl}\left(\overline{r}^{2};q\right)\right|\left(\mathop{\sum\sum}_{mn\equiv r(q)}\alpha_{m}\beta_{n}\gamma_{n,q}-\frac{1}{\phi(q)}\mathop{\sum\sum}_{(mn,q)=1}\alpha_{m}\beta_{n}\gamma_{n,q}\right),$
where $\boldsymbol{\beta}=(\beta_{n})$, $\boldsymbol{\gamma}_{q}=(\gamma_{n,q})$ are complex coefficients with support on $[N,2N]$ with $\lVert\boldsymbol{\gamma}_{q}\rVert_{\infty}\leq(\tau(q)\log 2q)^{C}$ for some $C>0$.
Let $r\geq 1$ and $M\geq N$. For any $A>0$, there exists some constant $B=B(A,C)>0$, such that
$\sum_{q\leq Q}\mu^{2}(q)\tau(q)^{r}\left|\varepsilon(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}_{q};q)\right|\ll\lVert\boldsymbol{\alpha}\rVert\lVert\boldsymbol{\beta}\rVert Q(MN)^{\frac{1}{2}}(\log MN)^{-A}$
for $Q\leq MN(\log MN)^{-B}$, where the implied constant depends only on $A$ and $r$.
Proof.
See Lemma 4.1 in [18]. ∎
Lemma 3.4.
For sufficiently large prime $p$, we have
$\mathop{{\sum}^{*}}_{a(p)}\left|\text{Kl}\left(\overline{a}^{2};p\right)\right|=\frac{8}{3\pi}p+O(\sqrt{p}\log p).$
Proof.
See Lemma 3.3 in [18]. ∎
Lemma 3.5.
Let $g$ and $(\lambda_{d})$ be given above. Let $k$ be a fixed positive integer and $F$ has a zero of order at least $k$ at 0. Then
$\sum_{n}g\left(\frac{n}{X}\right)\mu^{2}(n)k^{\omega(n)}\left(\sum_{d\mid n}\lambda_{d}\right)^{2}=4\widetilde{g}(1)c(k,F)\frac{X}{\log X}(1+o(1)),$
where $c(k,F)$ is defined by
$c(k,F)=\sum_{j=1}^{k}\frac{1}{\Gamma(j)^{2}}\binom{k}{j}\int_{0}^{1}F^{(j)}(x)^{2}(1-x)^{j-1}\left(3+x\right)^{j-1}\text{d}x.$
Proof.
See Proposition 4.1 in [17]. ∎
Before presenting the following lemma, let’s first establish some definitions. Define a smooth function $J(x)$ and assume it has a Taylor expansion
$J(x):=\sum_{n\geq 0}\frac{a_{n}}{n!}x^{n}$
[…]
$\check{J}_{N}(s):=\sum_{n\geq 0}\frac{a_{n}}{(s\log N)^{n}}.$
Lemma 3.6.
Let $N>1$ be not an integer. For any coefficient $\boldsymbol{y}=\left(y_{n}\right)$ with $y_{n}\ll\tau(n)^{O(1)}(\log n)^{O(1)}$, we have
$\sum_{n\leq N}y_{n}J\left(\frac{\log(N/n)}{\log N}\right)=\frac{1}{2\pi i}\int_{2-i\infty}^{2+i\infty}\check{J}_{N}(s)Y(s)\frac{N^{s}}{s}\text{d}s,$
@@
Let $v$, $v_{1}$, $v_{2}$ be fixed positive integers, $m$ be fixed even positive integer, $M>1$ be any real number. Let $P(x)$ and $Q(x)$ be two smooth functions with zeros at 0 of orders at least $v_{1}+m/2$, $v_{2}+m/2$, respectively. Let $Z(s_{1},s_{2})$ be holomorphic in the right half plane containing a neighborhood of $(0,0)$ with
@@
$R=(1+o(1))C_{0}Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{\Gamma(v)m!}\int_{0}^{1}P^{(v_{1}+m/2)}(x)Q^{(v_{2}+m/2)}(x)(1-x)^{v-1}\text{d}x.$
@@
We follow the proof of Lemma A.2 in [17] and use Goldston-Motohashi-Pintz-Yıldırım method in [6]. Assume
@@
From the condition of Lemma 3.7, this can be expanded as
@@
$\times(\xi+1)^{v+k_{1}-v_{1}+k_{2}-v_{2}-i}(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-i}.$
We only need to consider the term with the quantity $(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-m}$, since the term containing $(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-m}$ has the highest power with respect to $M$. Substituting the expression above into (4), we derive
$\frac{C_{0}}{(v+k_{1}-v_{1}+k_{2}-v_{2})!}\binom{v+k_{1}-v_{1}+k_{2}-v_{2}}{m}$ $Z^{(m)}(0,0)(\log M)^{v+k_{1}-v_{1}+k_{2}-v_{2}-m}$
(5) $\times\frac{1}{2\pi i}\int_{C_{3}}\frac{(\xi+1)^{k_{1}-v_{1}+k_{2}-v_{2}-m}}{\xi^{k_{2}-v_{2}-m/2+1}}\text{d}\xi.$
@@
$\begin{cases}&k_{1}-v_{1}\geq\frac{m}{2}\\ &k_{2}-v_{2}\geq\frac{m}{2},\end{cases}\ \begin{cases}&k_{1}-v_{1}<\frac{m}{2}\\ &k_{2}-v_{2}<\frac{m}{2},\end{cases}\ \begin{cases}&k_{2}-v_{2}\geq\frac{m}{2}\\ &k_{1}-v_{1}<\frac{m}{2},\end{cases}\ \begin{cases}&k_{2}-v_{2}<\frac{m}{2}\\ &k_{1}-v_{1}\geq\frac{m}{2}.\end{cases}$
In the last three cases, the integral evaluates to zero. When $k_{1}-v_{1}<\frac{m}{2}$ and $k_{2}-v_{2}<\frac{m}{2}$, only one pole $\xi=-1$ of order $m-(k_{1}-v_{1})-(k_{2}-v_{2})$ is in the region $\left\{\xi:|\xi|<2\right\}$. Then Cauchy’s residue theorem implies that
@@
When $k_{1}-v_{1}<\frac{m}{2}$ and $k_{2}-v_{2}\geq\frac{m}{2}$, we divide this case into two subcases, namely
@@
Only one pole $\xi=0$ of order $k_{2}-v_{2}-m/2+1$ is in the region $\left\{\xi:|\xi|<2\right\}$. Thus Cauchy’s residue theorem implies that
@@
and two poles $\xi=-1$ and $\xi=0$ are in the region $\left\{\xi:|\xi|<2\right\}$. Surprisingly, in this case, we can obtain
@@
thus the integral evaluates to zero.
The last case can be handled similarly to the third one. For simplicity, it is advantageous to interchange the roles of $s_{1}$ and $s_{2}$ (i.e. let $|s_{1}|=2\rho$, $|s_{2}|=\rho$) in this analysis.
Combining the above results, only the case
@@
$(1+o(1))C_{0}Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{m!}$
$\ \ \ \ \ \ \ \times\sum_{k_{1}\geq v_{1}+m/2}\sum_{k_{2}\geq v_{2}+m/2}\frac{a_{k_{1}}b_{k_{2}}}{(v+k_{1}-v_{1}+k_{2}-v_{2}-m)!}\binom{k_{1}-v_{1}+k_{2}-v_{2}-m}{k_{2}-v_{2}-m/2}$
$=$ $(1+o(1))C_{0}Z^{(m)}(0,0)\frac{(\log M)^{v-v_{1}-v_{2}-m}}{m!\Gamma(v)}\int_{0}^{1}P^{(v_{1}+m/2)}(x)Q^{(v_{2}+m/2)}(x)(1-x)^{v-1}\text{d}x,$